OpenStudy (anonymous):

Devaughn's age is two times Sydney's age. The sum of their ages is 51 . What is Sydney's age?

OpenStudy (kidrah69):

Welcome to openstudy! Lets post one question at a time! Also try not to post the same question it seems you posted this one twice :)

jagr2713 (jagr2713):

whats the blue around his question

OpenStudy (kidrah69):

And the help to your question is right below this question :) @jagr2713 the person wants QH's help

OpenStudy (kidrah69):

@dan815 is one of them

jagr2713 (jagr2713):

@kidrah69 u saw my message

OpenStudy (kidrah69):

http://openstudy.com/study#/updates/551f6a21e4b0a67c2b902aa4 <---come here :) @anonas1105

OpenStudy (preetha):

Thanks Mira and Pooja for helping. Anona did you get an explanation?

OpenStudy (swamijatin2):

you just take sydney's age x

OpenStudy (swamijatin2):

this question saying that devaughn's is 2 times of S therefore age of D be 2x

OpenStudy (swamijatin2):

acoording to question there sum is 51 -> X + 2X=51

OpenStudy (swamijatin2):

solve this

OpenStudy (kidrah69):

they left....

OpenStudy (alexandervonhumboldt2):

well yeah new users have manner to leave without ansering anything

OpenStudy (anonymous):

So S = 51/3 and D = 2times 51/3 = is 102/3 Correct?

Nnesha (nnesha):

translate into math form  3 ways to solve for x and y elimination substitution and graphing if you want to do graphing method the solution is where both lines intersect and remember they are asking sydney's age